OpenStudy (anonymous):

i'm struggling with mental arithmetic..could you guys give me some short tips to do fast mental arithmetic?and thanks!

hartnn (hartnn):

see some @AravindG tutorials on vedic maths.

OpenStudy (sepeario):

http://tedmuller.us/Math/MentalArithmetic.htm

Parth (parthkohli):

What I do for addition is this:\[63 + 45 = 60 + 40 + 5 + 3 = 100 + 8 = \boxed{108}\]For multiplication,\[21 \times 23 = 21\times20+21\times3 =420+63=\boxed{483}\]For subtraction,\[145 - 87 = 145 - 80 - 7 = 65 - 7 = \boxed{58}\]

Parth (parthkohli):

For division:\[168\div 21 = (84 \div 21)\times 2 = 4\times 2 = 8\]

OpenStudy (theviper):

nice explanation @ParthKohli :)

Parth (parthkohli):

Though it's just tricks. You need to discover these, and believe it or not: The first three tricks were discovered by me... and I didn't look 'em up!

Parth (parthkohli):

Thanks @TheViper! =)

OpenStudy (theviper):

U deserved that :D

Parth (parthkohli):

Another thing that I do for division is this:\[558\div 8 = 279\div 4 = 278\div 4 + 1\div 4 = 69.5 + 0.25 = 69.75\]

Parth (parthkohli):

On the contrary, I think that there's is nothing wrong with being slow with mental arithmetic; so is thought by Stephen Wolfram.

Parth (parthkohli):

And Conrad Wolfram.

Parth (parthkohli):

http://www.youtube.com/watch?v=60OVlfAUPJg

OpenStudy (anonymous):

wow..thanks guys..thanks @parthKohli @TheViper @Sepeario @hartnn ..i dont kno wat to say..thanks...

OpenStudy (theviper):

:)

hartnn (hartnn):

21*23 = 22^2-1 = 484 -1 =483 anything *25 = anything*100/4

Parth (parthkohli):

No problem!

Parth (parthkohli):

Yeah, and... \(n \times 5 = 10n/2\).

OpenStudy (anonymous):

@hartnn sorry ii dint get u...

Parth (parthkohli):

@mathavraj: It's a method described in Class 8 NCERT.

OpenStudy (anonymous):

hm well i was not under ncert syllabus..

hartnn (hartnn):

21*23 = (22-1)(22+1)=22^2-1^2 = 484-1 25*36 = 3600/4 = 900

Parth (parthkohli):

\[21\times 23 = (22 + 1)(22 - 1) = 22^2 - 1 = 484 - 1 = \boxed{483}\]

hartnn (hartnn):

i have so many tricks to share....but not now.

Parth (parthkohli):

It's easy to multiply by any multiple of \(5\). To multiply by 5, put a zero then halve. 10: put a zero or shift the decimal to the right. 15: \(n \times 15 = n\times10 + n \times 5\) 20: multiply by 2, put a zero.

Parth (parthkohli):

And so on...

hartnn (hartnn):

look up multiplication by series of 1 multiplication by series of 9 they are pretty simple.

OpenStudy (anonymous):

i understand u @hartnn ...do it as u pls..

Parth (parthkohli):

Oh yeah, another thing which Hartnn reminded me of!\[9n = 10n - n\]\[8n = 10n - 2n\]. . .

Parth (parthkohli):

\[9\times62 = 62\times10 - 62 = 620 - 62 = \boxed{558}\]

hartnn (hartnn):

multiplication by 9, 99, 999 .....even 9999999 ...can be done in seconds, mentally

Parth (parthkohli):

Use the rules of algebra and apply those on numbers: Variables are just numbers in disguise!

OpenStudy (anonymous):

@hartnn how?

OpenStudy (anonymous):

parth yeah u r rit..

Parth (parthkohli):

@mathavraj \(999n = 1000n - n\), \(9999n = 10000n - n\)... :P

hartnn (hartnn):

there's an another shorter way....